OpenStudy (natsent06):

Identify the set of side lengths that would form a right triangle. A) 6, 10, and 8 B) 7, 4, and 5 C) 5, 7, and 11 D) 4, 9, and 12

OpenStudy (shinalcantara):

if it is a right triangle, then this must be satisfied \[c^2 = a^2 + b^2\] let's have choice A 6,10 and 8 10 is the hypotenuse or the c in the pythagorean theorem \[c = \sqrt{6^2 + 8^2}\] simplify. what'd be the value?